An elevator manufacturer determines the maximum force acting on ascending passengers should not exceed 110% of a standing passenger's weight on their feet . What would be the minimum time possible according to this manufacturer to rise 10 floors from rest if each floor is 5.00 m apart?

(A) 6.00

(B) 10.0

(C) 19.0

(D) 24.0


1
Expert's answer
2021-04-26T17:04:43-0400

Balancing Forces,

"N-mg=ma"


From question,

"N<=\dfrac{110}{100}mg"

Put this value in above equation,

"1.1mg-mg=ma"

"a=0.1g"

Taking acceleration due to gravity, "g=10\space m/s^2"

Total displacement, "s=10\times 5=50\space m"

Initial velocity, "u=0\space m/s"

From Newton's laws of motion,

"s=ut+\dfrac{1}{2}at^2"

"50=\dfrac{1}{2}\times0.1gt^2"

"t=\sqrt{\dfrac{2\times50}{0.1g}}"

"t=\sqrt{\dfrac{100}{0.1\times10}}=10\space s"